TIFFCP(1) TIFFCP(1) NAME tiffsplit−splitamulti-imageTIFFintosingle-imageTIFFfiles SYNOPSIS tiffsplitsrc.tif[prefix] DESCRIPTION tiffsplittakesamulti-directory(page)TIFFfileandcreatesoneormoresingle-directory(page) TIFFfilesfromit.Theoutputfilesaregivennamescreatedbyconcatenatingaprefix,alexically orderedsuffixintherange[aa-zz],thesuffix.tif(e.g.xaa.tif,xab.tif,xzz.tif).Ifaprefixisnot specifiedonthecommandline,thedefaultprefixofxisused. OPTIONS None. BUGS Onlyaselectsetof‘‘knowntags’’iscopiedwhenspliting. SEEALSO tiffcp(1),tiffinfo(1),libtiff(3) September26,1994 1